OpenStudy (anonymous):

Describe circumstances where you would add exponents and when you would multiply exponents.

OpenStudy (mertsj):

if you are multiplying factors and the bases are the same, you add the exponents. If you are raising an expression to a power, you multiply the exponents.

OpenStudy (anonymous):

\( \large a^m \times a^n = a^{m+n} \) and \( \large ( a^m)^n = a^{m\times n}\)